OpenStudy (anonymous):

i really need help with this question ive been stuck on it for a while ...the volume of a spherical balloon is increasing at a constant rate of 5in^3/s. at what rate is the radius of the balloon increasing when its surface has area 16π in^2

OpenStudy (anonymous):

what is the formula for the volume of a sphere? we need that to start

OpenStudy (anonymous):

once we have that, this will take two steps to solve

OpenStudy (anonymous):

4/3πr^2

OpenStudy (anonymous):

ok so we have \[V(r)=\frac{4}{3}\pi r^3\] \[V'(r)=4\pi r^2 r'\] you are told \(V'(r)=5\) solve for \(r'\)

OpenStudy (anonymous):

oh damn we also need \(r\) don't we. that is ok, we find the formula for the surface area and solve for \(r\) ok i lied it is 3 steps

OpenStudy (anonymous):

surface area is 4πr^2 right?

OpenStudy (anonymous):

\[S(r)=4\pi r^2\] \[16\pi =4\pi r^2\] \[4=r^2\] \[r=2\] now we have \(r\)

OpenStudy (anonymous):

and we have \[5=4\pi r^2r'\] replace \(r\) by 2 get \[5=16\pi r'\] solve for \(r'\)

OpenStudy (anonymous):

do you just minus 16 from both sides then minus π? im very confused at this point

OpenStudy (anonymous):

how do i solve 5=16πr' for r'?